OpenStudy (anonymous):

Find the y-coordinate for the solution of the system of linear equations. 5x - 3y = -11 -7x - y = 5

myininaya (myininaya):

Ok the easiest way for me to solve this would be to multiply the bottom equation by -3 and then add the equations

myininaya (myininaya):

5x-3y=-11 -> 5x-3y=-11 -7x-y=5 times -3-> 21x+3y=-15 --------------- add 26x+0=-26 26x=-26

myininaya (myininaya):

Does that look as pretty as it does to me on my screen?

OpenStudy (anonymous):

Yes it does

myininaya (myininaya):

Great so what is x?

OpenStudy (anonymous):

26

myininaya (myininaya):

Try again We have 26x=-26

OpenStudy (anonymous):

-26 lol

myininaya (myininaya):

26x = -26 -- --- Divided both sides by 26 26 26

myininaya (myininaya):

26/26=1 and -26/26=-1 so we have 26x = -26 -- ---- 26 26 => 1x = -1 => x = -1

OpenStudy (anonymous):

oh ok i got it now thanx

myininaya (myininaya):

But we want to know what y is... So take this -1 and go to one of the equations that were given to you and replace x with -1 and solve for y and then you will be done :)

OpenStudy (apoorvk):

great patience and explanation work @myininaya. you inspire!!

OpenStudy (anonymous):

so the answer would be 1

myininaya (myininaya):

So this is one of the equations given to you: 5x - 3y = -11 and we found x to be -1 So replacing x with -1 gives me 5(-1)-3y=-11 -5-3y=-11 Can you solve this for y?

OpenStudy (anonymous):

no i cant

myininaya (myininaya):

-5 - 3y = -11 To solve for y (or to isolate y) some possible steps to take are: To first add 5 on both sides -5 - 3y = -11 +5 +5 --------------------- 0 - 3y = -6 -3y = -6 Can you solve this equation for y? I know you can do it :) Try.

OpenStudy (anonymous):

it would be 2